A girl is half her mothers age..the father is 5 years olfer than the mother..when we add all theit ages it gives us 100...how old is the mother , girl and the father
Answer by Cromlix(4381)

Mother's age = x
Girl's age = x/2
Father's age = x + 5
x/2 + x + (x + 5) = 100
x/2 + x + x + 5 = 100
Common denominator = 2
x + 2x + 2x + 10 = 200
5x + 10 = 200
5x = 190
x = 38
Mother's age = 38 years old
Girl's age = 19 years old
Father's age = 43 years old.
Hope this helps.
:-)